# Diffusion-Driven Generative Framework for Molecular Conformation Prediction

### 22 Dec 2023 v1

@@ IV-B Reverse Process
Our objective is to infer the original molecular conformations, denoted by $\mathbf{c}^{0}$, starting from a state of maximum entropy represented by white noise $\mathbf{c}^{T}$, conditioned on the provided molecular graph structures $G$. We approach this inference task as the temporal reversal of the diffusion sequence, initiating from the high-entropy state where the conformational coordinates $\mathbf{c}^{T}$ are distributed according to $p(\mathbf{c}^{T})$. We formulate this reverse dynamics as a conditional Markov chain with learnable transitions:
@@
In our framework, neural networks parameterized by $\theta$ are employed to approximate mean values $\mu_{\theta}$, while $\sigma_{t}$ represents a variance that can be arbitrarily chosen by the practitioner. The process commences with the assignment of the initial state distribution $p(\mathbf{c}^{T})$ as a standard Gaussian distribution. For any given molecular graph $G$, its three-dimensional configuration is constructed by first sampling disordered state particles $\mathbf{c}^{T}$ from $p(\mathbf{c}^{T})$, which are subsequently progressively sculpted into the desired structure via a sequence of reverse Markov transitions $p_{\theta}(\mathbf{c}^{t-1}|G,\mathbf{c}^{t})$.
@@ IV-C Training Objective
In order to account for the multiplicity of stable molecular structures, our study implements a conditional variational autoencoder (CVAE) framework to construct the probabilistic model $p(\mathbf{C}|G)$. Within this framework, we introduce a latent variable $z$ to encapsulate the intrinsic uncertainties associated with the generation of molecular conformations. The architecture of the CVAE comprises a prior latent distribution $p_{\psi}(z|G)$ along with a decoding mechanism $p_{\theta}(\mathbf{C}|z,G)$ which delineates the conditional likelihood of the conformation $\mathbf{C}$ for a given $z$. Concurrently, the training phase is augmented with an inferential component (encoder) $q_{\psi}(z|\mathbf{C},G)$, enhancing the model’s ability to learn from data. The encoder and decoder are jointly trained to maximize the evidence lower bound (ELBO) of the data log-likelihood:
@@
The ELBO, fundamental to our model’s training regime, is essentially the sum of two distinct components: the negative reconstruction loss, $L_{recon}$, and the latent space regularization term, $L_{prior}$. The former measures the fidelity of reconstructed data to the original, while the latter ensures the latent variable distribution remains well-behaved, or close to the prior. To encapsulate the model’s probabilistic nature, both the encoder’s output $q_{\phi}(z|\mathbf{C},G)$ and the prior distribution $p_{\psi}(z|G)$ are formulated as diagonal Gaussian distributions $\mathcal{N}(z|\mu_{\phi}(\mathbf{C},G),\sigma_{\phi}(\mathbf{C},G))$ and $\mathcal{N}(z|\mu_{\psi}(G),\sigma_{\psi}(G))$, respectively. The parameters of these distributions, namely the means and standard deviations, are deduced by specialized graph neural networks. In the process of optimizing the ELBO, a reparametrization trick is employed to enable backpropagation through the stochastic sampling process. Specifically, a latent variable $z$ is sampled by first calculating $\mu_{\phi}(\mathbf{C},G)$ and $\sigma_{\phi}(\mathbf{C},G)$, then adding a scaled random noise component $\epsilon$, where $\epsilon$ follows a standard Gaussian distribution $\mathcal{N}(0,I)$. This methodology facilitates gradient-based optimization by making the sampling process differentiable.
@@ IV-D Inner Objective
To circumvent the rotational and translational dependency inherent in Cartesian coordinate generation for molecular conformations, modern methods [17] typically employ an alternate strategy. These methods utilize a decoder architecture designed to produce inter-atomic distance matrices, with $d_{\theta,\phi}=D_{\theta}(z_{\phi},G)$ representing the decoded distances conditioned on the latent variable $z$ and the graph $G$. These distance matrices serve as a translationally and rotationally invariant representation of the molecule’s geometry. After the distance matrix $d$ is generated, it can then be used as a basis for reconstructing the three-dimensional conformation $\mathbf{C}$. This reconstruction is achieved by addressing a distance geometry problem, which essentially entails finding a set of coordinates that best comply with the given inter-atomic distances. The optimization process involved is often a complex numerical problem that requires sophisticated techniques to resolve the coordinates in a way that faithfully respects the distances while also conforming to physical plausibility.
@@ IV-E Outer Objective
Our primary aim is to refine the generation of 3D molecular structures, aligning them as closely as possible to the actual models, regardless of their orientation or positioning. To measure this alignment’s precision, we employ the post-alignment Root-Mean-Square Deviation (RMSD), a standard gauge in the field. This metric is determined after optimally aligning a generated conformation $\mathbf{C}$ to a reference $\mathbf{C}^{*}$ using an alignment operation $A$, ensuring minimal disparity between the two. The RMSD is thus defined as the square root of the average squared distance between the atoms of the aligned generated conformation and the reference after this operation is applied.
@@ IV-F Bilevel Program
To navigate the complexities of molecular conformation generation, our approach adopts a nested optimization scheme, where we distinguish between two intertwined objectives: one at a macro-level and another at a micro-level. The macro-level objective (denoted as the outer problem in Equation 10) is to capture the conditional distribution $p(\mathbf{C}|G)$ of molecular structures, encapsulating the full breadth of conformations that molecules can adopt given their underlying graph $G$. Concurrently, the micro-level objective (referred to as the inner problem in Equation 8) is dedicated to determining the precise molecular conformation from a collection of interatomic distances that are predicted. By taking the expectation over latent variable $z$, the resulting bilevel program for calculating the reconstruction term $L_{reconin}$ equation 7 can be written as:
@@
Addressing the complexity of the bilevel optimization in our model, we circumvent the lack of closed-form solutions and intractable expectations by leveraging a variational inference approach. This is paired with the reparametrization trick to provide a tractable, gradient-based estimation of molecular conformations, a technique we detail in the subsequent sections.
@@ IV-G Generative Model
Our generative model for molecular conformations is underpinned by the use of message-passing neural networks (MPNNs) [45], a class of graph neural networks renowned for their exemplary representation learning capabilities for molecular structures. MPNNs work on the principle of updating atom embeddings through an aggregation of neighboring node information within the graph structure $G$, ensuring invariance to graph isomorphism and facilitating state-of-the-art molecular modeling [46, 47, 48]. We have adopted the MPNN architecture akin to the one delineated [17] for the encoder $q_{\phi}(z|\mathbf{C},G)$ and the prior $p_{\psi}(z|G)$. Recognizing the intensive memory demands of bilevel optimization, our decoder $p_{\theta}(\mathbf{C}|z,G)$ is innovatively implemented using a continuous normalizing flow (CNF) based on an ordinary differential equation (ODE), offering the advantage of a fixed memory footprint regardless of the number of layers. Details of our decoder’s design are provided in the subsequent sections.
In our design, the decoder operates through two sequential stages. The initial stage involves a distance prediction network $D_{\theta}(z,G)$ which interprets the latent space variable $z$ into an array of interatomic distances, denoted by $d$. Subsequently, the model employs a differential geometry algorithm that is capable of reconstructing the molecular conformation $\mathbf{C}$ from the set of distances $d$. The network $D_{\theta}(z,G)$ is a conditional adaptation of the CNF model which commences with a noise distribution $d(t_{0})$ representing the starting distances in the CNF ODE path originating from a standard Gaussian distribution $\mathcal{N}(0,I)$ and culminates with the ultimate distances $d=d(t_{1})$. The transformation is conditioned on the latent variable $z$ as well as the graph $G$:
@@
In the training of our molecular conformation generative model, we employ an ODE solver to integrate the continuous-time dynamics for backpropagation, complemented by an auxiliary objective $L_{aux}$ that imposes a regularizing effect on distance predictions. This integration ensures that the model’s parameters are optimized to yield accurate molecular geometries, harnessing the ODE solver’s gradient estimations and the guiding influence of $L_{aux}$ to refine the distance predictions crucial for generating realistic molecular conformations. In summary, the training objective can be interpreted as the sum of three parts:
@@ IV-H Bilevel Optimization
To tackle the optimization of the bilevel problem articulated in equation 11 and equation 12, we propose a pragmatic algorithm that builds on existing methods for inferring 3D structures from interatomic distances. The inner problem, represented by equation 12, is akin to traditional distance geometry challenges commonly encountered in molecular structure determination. Analogous problems have been approached through semi-definite programming (SDP) for deducing protein configurations from NMR data, as well as employing an Alternating Direction Method of Multipliers (ADMM) to computationally fold proteins into their native three-dimensional forms. Drawing inspiration from these methodologies, our algorithm is designed to effectively navigate the complexity of the inner optimization problem, thereby facilitating accurate reconstruction of molecular geometries from predicted interatomic distances. This approach allows us to efficiently generate 3D coordinates that are consistent with the given pairwise distances, which is critical for the successful application of our model to the prediction and generation of molecular conformations.
@@
In the process of model training, we concentrate on determining the hypergradient by computing the derivative of the expected reconstruction error with respect to the parameters, which influences the adjustment of the model to enhance conformation predictions. This is achieved by generating molecular conformations using a series of gradient descent steps, signified by $\mathbf{C}_{\theta,\phi,T}$, to approximate the solution for the distance geometry problem.
@@
To compute the gradient $\nabla_{\theta,\phi}\mathbf{C}_{\theta,\phi,T}$, we fully trace the evolution of the inner optimization from the end state $\mathbf{c}_{T}$ back to the initial state $\mathbf{c}_{0}$. During the forward pass, we store each intermediary geometry $\mathbf{c}_{0,\cdots,T}$ resulting from the inner optimization dynamics. For the backward computation, these saved geometries facilitate the calculation of gradients via a sequence of Vector-Jacobian Products (VJPs). During the reverse computation, the gradient starting from the $\partial_{\mathbf{c}_{T}}F$ can be propagated to the intermediate geometries $\mathbf{c}_{t}$ through $\nabla_{\mathbf{c}_{t}}\mathbf{c}_{t+1}$:
@@
IV-I Sampling
To synthesize a molecular conformation $\mathbf{C}$ from a given molecular graph $G$, we first sample a latent variable $\tilde{z}$ from the learned prior distribution $p_{\psi}(z|G)$. Next, we obtain a set of initial interatomic distances $d(t_{0})$ by sampling from a Gaussian distribution. These sampled distances $\tilde{d}(t_{0})$ are processed by the trained Neural ODE $G_{\theta}$, which, given the conditions $\tilde{z}$ and $G$, transforms $\tilde{d}(t_{0})$ into a new set of distances $\tilde{d}$. Finally, we use these predicted distances $\tilde{d}$ to reconstruct the molecular conformation $\mathbf{C}$ by solving the optimization problem $\text{arg min}_{\mathbf{C}}H(\mathbf{C},d_{\theta,\phi})$, which aims to find the conformation $R$ that best matches the distances, as defined in the previously referred equation.
@@ IV-I Sampling
1:  Sample $\mathbf{c}^{T}\sim p(\mathbf{c}^{T})=\mathcal{N}(0,I)$
@@
3:     Compute $\mu_{\theta}(\mathbf{c}^{s},G,s)$ from $\epsilon_{\theta}(\mathbf{c}^{s},G,s)$;
4:     Sample $\mathbf{c}^{s-1}\sim\mathcal{N}(\mathbf{c}^{s-1};\mu_{\theta}(\mathbf{c}^{s},G,s),\sigma^{2}_{t}I)$
5:     Calculate $\mu_{q},\sigma_{q}$;
6:     Calculate $\mathcal{L}_{prior}$;
7:     $d^{*}\leftarrow R_{t}^{*}$;
8:     Calculate $\mathcal{L}_{aux}$;
9:     Initialize $\mathbf{0}^{0}$, sample $d(t_{0})\sim\mathcal{N}(0,I)$;
@@
11:     for $t=1,2,\cdots,T$ do
12:        $\mathbf{c}^{t+1}=\mathbf{c}^{t}-\eta\nabla H(\mathbf{c}^{t},d)$;
13:     end for
14:     $R\leftarrow\mathbf{c}^{T}$;
15:     Calculate $\mathcal{L}_{recon}$;
16:     Calculate $\mathcal{L}$;
17:     Update $\theta,\phi,\psi$;
18:  end for
19:  return  $q_{\phi},p_{\theta},p_{\psi}$

### 21 Jan 2024 v2

@@ IV-B Reverse Process
Our goal is to deduce the initial molecular conformations, denoted as $\mathbf{c}^{0}$, starting from a state of maximum entropy represented by white noise $\mathbf{c}^{T}$. This process is conditioned on the given molecular graph structures denoted as $G$. We approach this inference task as the temporal reversal of the diffusion sequence, commencing from the high-entropy state where the conformational coordinates $\mathbf{c}^{T}$ follow the distribution $p(\mathbf{c}^{T})$. We formulate this reverse dynamics as a conditional Markov chain with learnable transitions:
@@
In our framework, we utilize neural networks parameterized by $\theta$ to approximate mean values denoted as $\mu_{\theta}$, while $\sigma_{t}$ represents a variance that can be chosen by the practitioner. The process begins with the assignment of the initial state distribution $p(\mathbf{c}^{T})$ as a standard Gaussian distribution. For any given molecular graph $G$, its three-dimensional configuration is constructed by initially sampling disordered state particles $\mathbf{c}^{T}$ from $p(\mathbf{c}^{T})$. These particles are then progressively molded into the desired structure through a sequence of reverse Markov transitions $p_{\theta}(\mathbf{c}^{t-1}|G,\mathbf{c}^{t})$.
@@ IV-C Training Objective
To account for the multiplicity of stable molecular structures, our study employs a conditional variational autoencoder (CVAE) framework to construct the probabilistic model $p(\mathbf{C}|G)$. Within this framework, we introduce a latent variable $z$ to capture the intrinsic uncertainties associated with the generation of molecular conformations. The CVAE architecture consists of a prior latent distribution $p_{\psi}(z|G)$ and a decoding mechanism $p_{\theta}(\mathbf{C}|z,G)$, which defines the conditional likelihood of the conformation $\mathbf{C}$ given $z$. Simultaneously, the training process incorporates an inferential component (encoder) $q_{\psi}(z|\mathbf{C},G)$, enhancing the model’s ability to learn from data. The encoder and decoder are jointly trained to maximize the evidence lower bound (ELBO) of the data log-likelihood:
@@
This framework allows us to probabilistically model molecular conformations while accounting for inherent uncertainties and dependencies within the data. The Evidence Lower Bound (ELBO), central to our model’s training regimen, consists of two distinct components: the negative reconstruction loss, denoted as $L_{recon}$, and the latent space regularization term, represented as $L_{prior}$. The former assesses how well the reconstructed data matches the original data, while the latter promotes the maintenance of a well-behaved latent variable distribution, approximating the prior. To capture the probabilistic nature of the model, both the output of the encoder, $q_{\phi}(z|\mathbf{C},G)$, and the prior distribution, $p_{\psi}(z|G)$, are defined as diagonal Gaussian distributions, with parameters expressed as $\mathcal{N}(z|\mu_{\phi}(\mathbf{C},G),\sigma_{\phi}(\mathbf{C},G))$ and $\mathcal{N}(z|\mu_{\psi}(G),\sigma_{\psi}(G))$, respectively. The means and standard deviations of these distributions are determined by specialized graph neural networks. During the optimization of the ELBO, we employ a reparametrization trick to enable backpropagation through the stochastic sampling process. Specifically, the latent variable $z$ is sampled by first computing $\mu_{\phi}(\mathbf{C},G)$ and $\sigma_{\phi}(\mathbf{C},G)$, and then adding a scaled random noise component $\epsilon$, where $\epsilon$ follows a standard Gaussian distribution $\mathcal{N}(0,I)$. This methodology facilitates gradient-based optimization by rendering the sampling process differentiable.
@@ IV-D Inner Objective
To mitigate the challenges associated with rotational and translational dependencies in the generation of Cartesian coordinates for molecular conformations, contemporary methods often adopt an alternative approach. These techniques employ a decoder architecture specifically designed to generate inter-atomic distance matrices, denoted as $d_{\theta,\phi}=D_{\theta}(z_{\phi},G)$, where these decoded distances are conditioned on the latent variable $z$ and the graph $G$. These distance matrices serve as a representation of the molecule’s geometry that is invariant to translation and rotation. Once the distance matrix $d$ is generated, it serves as the foundation for reconstructing the three-dimensional conformation $\mathbf{C}$. This reconstruction process involves solving a distance geometry problem, which essentially entails finding a set of coordinates that best align with the provided inter-atomic distances. The optimization process for this problem is often intricate and requires sophisticated techniques to determine the coordinates in a manner that accurately reflects the distances while adhering to physical plausibility.
@@ IV-E Outer Objective
Our primary objective is to enhance the generation of 3D molecular structures in a way that closely aligns them with the actual models, regardless of their orientation or positioning. To quantify the precision of this alignment, we utilize the post-alignment Root-Mean-Square Deviation (RMSD), a commonly used metric in the field. This metric is calculated after optimally aligning a generated conformation $\mathbf{C}$ with a reference conformation $\mathbf{C}^{*}$ using an alignment operation $A$ that minimizes the disparity between the two. The RMSD is defined as the square root of the average squared distance between the atoms of the aligned generated conformation and the reference conformation after the alignment operation has been applied.
@@ IV-F Bilevel Program
To tackle the intricacies of molecular conformation generation, our approach employs a nested optimization scheme, where we distinguish between two interrelated objectives: one at a macro-level and another at a micro-level. The macro-level objective (referred to as the outer problem in Equation 10) aims to capture the conditional distribution $p(\mathbf{C}|G)$ of molecular structures, encompassing the full spectrum of conformations that molecules can assume based on their underlying graph $G$. Simultaneously, the micro-level objective (referred to as the inner problem in Equation 8) is focused on determining the precise molecular conformation from a set of predicted interatomic distances. By taking the expectation over the latent variable $z$, the resulting bilevel program for calculating the reconstruction term $L_{reconin}$ in Equation 7 can be expressed as follows:
@@
To handle the complexity of the bilevel optimization in our model, we address the absence of closed-form solutions and intractable expectations by utilizing a variational inference approach. This, combined with the reparametrization trick, offers a tractable, gradient-based estimation of molecular conformations, a technique we elaborate on in the following sections.
@@ IV-G Generative Model
Our generative model for molecular conformations relies on the use of message-passing neural networks (MPNNs) [47], a class of graph neural networks known for their exceptional representation learning capabilities for molecular structures. MPNNs operate by updating atom embeddings through the aggregation of neighboring node information within the graph structure $G$, ensuring invariance to graph isomorphism and enabling state-of-the-art molecular modeling [48, 49, 50]. We have employed an MPNN architecture similar to the one described in [18] for the encoder $q_{\phi}(z|\mathbf{C},G)$ and the prior $p_{\psi}(z|G)$. Recognizing the intensive memory requirements of bilevel optimization, our decoder $p_{\theta}(\mathbf{C}|z,G)$ is innovatively implemented using a continuous normalizing flow (CNF) based on an ordinary differential equation (ODE), offering the advantage of a fixed memory footprint regardless of the number of layers. Details of our decoder’s design are provided in the following sections.
In our architectural design, the decoder operates through two sequential stages. The initial stage involves a distance prediction network $D_{\theta}(z,G)$, which translates the latent space variable $z$ into an array of interatomic distances denoted by $d$. Subsequently, the model utilizes a differential geometry algorithm capable of reconstructing the molecular conformation $\mathbf{C}$ from the set of distances $d$. The network $D_{\theta}(z,G)$ is a conditional adaptation of the continuous normalizing flow (CNF) model. It begins with a noise distribution $d(t_{0})$, representing the initial distances in the CNF ordinary differential equation (ODE) path, originating from a standard Gaussian distribution $\mathcal{N}(0,I)$. The path concludes with the final distances $d=d(t_{1})$. This transformation is conditioned on the latent variable $z$ as well as the graph $G$:
@@
In the training of our molecular conformation generative model, we utilize an ODE solver to integrate the continuous-time dynamics for backpropagation. This is complemented by an auxiliary objective $L_{aux}$, which imposes regularization on distance predictions. The integration process ensures that the model’s parameters are optimized to produce accurate molecular geometries. It leverages the gradient estimations from the ODE solver and the guidance provided by $L_{aux}$ to enhance the accuracy of distance predictions, a crucial aspect of generating realistic molecular conformations. In summary, the training objective can be considered as the sum of three components:
@@ IV-H Bilevel Optimization
To address the optimization of the bilevel problem described in equation 11 and equation 12, we propose a practical algorithm that builds upon existing methods used for inferring 3D structures from interatomic distances. The inner problem, as expressed in equation 12, is analogous to traditional distance geometry challenges commonly encountered in molecular structure determination. Similar problems have been tackled using techniques such as semi-definite programming (SDP) for deducing protein configurations from NMR data and employing the Alternating Direction Method of Multipliers (ADMM) to computationally fold proteins into their native three-dimensional forms. Taking inspiration from these methodologies, our algorithm is designed to effectively address the intricacies of the inner optimization problem, facilitating the accurate reconstruction of molecular geometries from predicted interatomic distances. This approach enables the efficient generation of 3D coordinates that align with the provided pairwise distances, a critical aspect for the successful application of our model in predicting and generating molecular conformations.
@@
During the model training process, our focus is on calculating the hypergradient, which involves computing the derivative of the expected reconstruction error with respect to the model parameters. This derivative guides the adjustment of the model to improve conformation predictions. To compute the hypergradient, we generate molecular conformations through a series of gradient descent steps denoted as $\mathbf{C}_{\theta,\phi,T}$, which approximates the solution for the distance geometry problem.
@@
To compute the gradient $\nabla_{\theta,\phi}\mathbf{C}_{\theta,\phi,T}$, we meticulously trace the entire evolution of the inner optimization process, starting from the end state $\mathbf{c}_{T}$ and working our way back to the initial state $\mathbf{c}_{0}$. During the forward pass, we store each intermediate geometry $\mathbf{c}_{0,\cdots,T}$ that arises from the inner optimization dynamics. When performing the backward computation, these saved geometries play a crucial role in calculating gradients through a sequence of Vector-Jacobian Products (VJPs). During the reverse computation, we propagate the gradient starting from $\partial_{\mathbf{c}_{T}}F$ to the intermediate geometries $\mathbf{c}_{t}$ using $\nabla_{\mathbf{c}_{t}}\mathbf{c}_{t+1}$:
@@
1:  Sample an initial molecular conformation $\mathbf{c}^{T}$ from a standard Gaussian distribution $\mathcal{N}(0,I)$
@@
3:     Calculate $\mu_{\theta}(\mathbf{c}^{s},G,s)$ from $\epsilon_{\theta}(\mathbf{c}^{s},G,s)$;
4:     Sample the next conformation $\mathbf{c}^{s-1}$ from a Gaussian distribution with mean $\mu_{\theta}(\mathbf{c}^{s},G,s)$ and variance $\sigma^{2}_{t}I$
5:     Compute $\mu_{q}$ and $\sigma_{q}$ and calculate the $\mathcal{L}_{prior}$ term.
6:     $d^{*}\leftarrow R_{t}^{*}$;
7:     Initialize the distance variables as $d^{0}$ and sample the initial distances $d(t_{0})$ from a Gaussian distribution;
@@
9:     for $t=1,2,\cdots,T$ do
10:        $\mathbf{c}^{t+1}=\mathbf{c}^{t}-\eta\nabla H(\mathbf{c}^{t},d)$;
11:     end for
12:     $R\leftarrow\mathbf{c}^{T}$;
13:     Calculate the reconstruction loss $\mathcal{L}_{recon}$;
14:     Calculate the total loss $\mathcal{L}$;
15:     Update $\theta,\phi,\psi$;
16:  end for
17:  return  $q_{\phi},p_{\theta},p_{\psi}$
IV-I Sampling
To generate a molecular conformation $\mathbf{C}$ based on a given molecular graph $G$, our process unfolds as follows: Firstly, we begin by sampling a latent variable $\tilde{z}$ from the learned prior distribution $p_{\psi}(z|G)$. Next, we acquire an initial set of interatomic distances $d(t_{0})$ by sampling from a Gaussian distribution. These sampled distances, denoted as $\tilde{d}(t_{0})$, are a starting point for the distance prediction. These sampled distances $\tilde{d}(t_{0})$ undergo processing through the trained Neural ODE $G_{\theta}$. This network, conditioned on $\tilde{z}$ and $G$, transforms the initial distances $\tilde{d}(t_{0})$ into a new set of distances, denoted as $\tilde{d}$. Finally, we utilize these predicted distances $\tilde{d}$ to reconstruct the molecular conformation $\mathbf{C}$. This is achieved by solving the optimization problem $\text{arg min}_{\mathbf{C}}H(\mathbf{C},d_{\theta,\phi})$, which seeks to find the conformation $R$ that best matches the predicted distances, as defined in the previously mentioned equation.